If sodium-24 has a half-life of fifteen hours, and you have a sample that contains 15 kg of sodium-24, how much would remain after twelve hours?
Answer by stanbon(75887)

If sodium-24 has a half-life of fifteen hours, and you have a sample that contains 15 kg of sodium-24, how much would remain after twelve hours?
----
A(t) = Ao(1/2)^(t/k) where k is the half-life
-----------------------------
A(12) = 15(1/2)^(12/15)
= 15*0.57435
---
= 8.6152 kg
-------------
=====
Cheers,
Stan H.